OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. 3/x + 3/x+ -3/x+2

OpenStudy (bahrom7893):

Find the common denominator by multiplying x by x+2. The new common denominator will be x(x+2), and the resulting fraction is going to be: [3(x+2)+3(x+2)-3x]/[x(x+2)]

OpenStudy (anonymous):

are you sure your reading the problem right

OpenStudy (anonymous):

its 3/x + 3/x- 3/x+2

OpenStudy (bahrom7893):

Can you draw it or use parenthesis properly?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

can i send you a picture

OpenStudy (anonymous):

of it

OpenStudy (bahrom7893):

yea attach a picture

OpenStudy (anonymous):

ill just send you link

OpenStudy (anonymous):

ok ill attach it give me a secx

OpenStudy (anonymous):

u there

OpenStudy (bahrom7893):

Ok just post the correct question next time.. You wrote 3/x + 3/x- 3/x+2 It's actually 3/x + 3/(x+1) + 3/(x+2)

OpenStudy (anonymous):

ok sry about that

OpenStudy (bahrom7893):

So find the common denominator by multiplying all the denominators. The common denominator will be: x*(x-1)*(x+2)

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

Sorry i just realized that the question said: 3/x + 3/(x+1) - 3/(x+2). The first step will remain the same though. Combine all the fractions under one common denominator: [3(x+1)(x+2)+3x(x+2)-3x(x+1)]/[x(x+1)(x+2)]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok whats the restrictions

OpenStudy (bahrom7893):

Now simplify the numerator by expanding all the parenthesis: [3(x+1)(x+2)+3x(x+2)-3x(x+1)]/[x(x+1)(x+2)] = [3(x^2+2x+x+2x+x^2+2x-x^2-x)]/[x(x+1)(x+2)] = [3(x^2+6x)]/[x(x+1)(x+2)]= 3x(x+6)/[x(x+1)(x+2)]

OpenStudy (anonymous):

is this step 2

OpenStudy (bahrom7893):

it's step 2 continued.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

continue plz

OpenStudy (bahrom7893):

And finally simplify the fraction further by cancelling Xs in the numerator and the denominator to get: (3x+18)/[(x+1)(x+2)]

OpenStudy (anonymous):

ok whats the restrictions

OpenStudy (bahrom7893):

The restrictions are as follows: x cannot be negative one or negative two.

OpenStudy (anonymous):

ok ty so much

OpenStudy (anonymous):

can you heelp me with another 1

OpenStudy (anonymous):

its easier